# Understanding Iterative Combinatorial Auction Designs via Multi-Agent Reinforcement Learning

### v1

@@ 5.3. Theoretical Questions
In two-player zero-sum games, CFR is guaranteed to converge to an approximate Nash equilibrium. However, auctions are not zero-sum, and existing guarantees for general-sum games are far weaker. Morrill et al. (2021) show that CFR is only guaranteed to converge to a counterfactual coarse correlated equilibrum, a much weaker solution concept. Despite this lack of theoretical justification, on our auctions, we found that MCCFR consistently converged to pure Bayes–Nash (near-)equilibrium. Is there a class of games (larger than two-player zero-sum) in which existing MARL algorithms converge to Nash equilibria? Understanding the space of such games would help with selecting an appropriate MARL algorithm.
@@
Acknowledgements
Thanks to Marc Lanctot and Betty Shea for their technical contributions on early iterations of this work, and to Costa Huang, Paul Milgrom, Julien Perolat, Samuel Sokota, and Albert Zuo for helpful discussions. This work was supported in part by an NSERC CGS-D scholarship, and through computational resources and services provided by Advanced Research Computing at the University of British Columbia. This work was done in part while several authors were visiting the Simons Institute for the Theory of Computing.
@@ Appendix A Computational Environment
Our experiments were performed on two different compute clusters. Our first cluster consisted of nodes equipped with 32 2.10GHz Intel Xeon E5-2683 v4 CPUs with 40960 KB cache and 96 GB RAM. Our second cluster consisted of 16-core machines with Intel Silver 4216 Cascade Lake 2.1GHz processors and 96 GB RAM. Jobs were scheduled such that each run had access to 4 CPUs and 20 GB of RAM.
@@ Appendix C Value Sampling Details
Our game sampler rejects games that we deem strategically uninteresting. We require that no bidder, in all type realizations, is allocated nothing under the utility-maximizing allocation at opening prices. We solve for the utility-maximizing allocation using a MIP. Such a bidder is hopelessly weak, tending to slow down simulations for no added benefit.
@@
Some criteria ensured that we generated games that we were capable of solving. First, as a proxy measure for the length of the auction, we computed the number of rounds it would take for the auction to conclude in a simulation in which each bidder bids for its profit-maximizing bundle at start-of-round prices in every round. We rejected auctions longer than 20 such rounds. Secondly, we reran the above simulation, this time incrementing the price on a single, randomly chosen product among those that were overdemanded in each round. This is a crude model of the worst-case auction (from a length perspective), in which each round prices only increment on a single product. We rejected simulations that took on average longer than 25 such rounds to conclude. Lastly for 2-player games, we required MCCFR to be able to run at least 25 iterations per second. These checks were all based on the drop-by-player rules.
@@
We made several small changes to the value model for the 3-player games in Section 4.3. For SATS parameters, bidders had market shares sampled uniformly from 20% to 30% and values per subscriber sampled uniformly from \$35 million to \$45 million. The keypoints on their sigmoidal value function were placed at $\pm$10from their market share. We also adjusted the rejection sampler requirement that MCCFR be able to run 25 iterations per second, loosening this threshold to 1 iteration per second.
@@
Appendix D PPO Hyperparameters
@@
        optimizer_kwargs[’betas’] = [beta1, beta2]
    config[’optimizer_kwargs’] = optimizer_kwargs
@@
    agent_kwargs = dict()
@@
        agent_kwargs[’activation’] = str(np.random.choice([’relu’, ’tanh’]))
        agent_kwargs[’hidden_sizes’] = random.choice([
            [32, 32, 32], [64, 64, 64], [128, 128, 128],
        agent_kwargs[’add_skip_connections’] = bool(np.random.choice([True, False]))
        agent_kwargs[’use_torso’] = bool(np.random.choice([True, False]))
@@
        agent_kwargs[’actor_hidden_sizes’] = hidden_sizes
        agent_kwargs[’critic_hidden_sizes’] = hidden_sizes
        agent_kwargs[’actor_activation’] = activation
        agent_kwargs[’critic_activation’] = activation
    config[’agent_fn_kwargs’] = agent_kwargs
    return config
Our final PPO configuration is as follows:
@@
0 Activation relu
1 Add Skip Connections True
2 Anneal LR True
3 Architecture AuctionNet
4 Clip Coef 0.018894
5 Clip Vloss False
6 Entropy Coef 0.000030
7 Gae False
8 Gae Lambda 0.985872
9 Gamma 1.000000
10 Hidden Sizes [32, 32, 32]
11 Learning Rate 0.000090
12 Max Grad Norm 0.526474
13 Normalize Advantages False
14 Num Annealing Updates
15 Num Envs 16
16 Num Minibatches 8
17 Optimizer adam
18 Optimizer Betas [0.8032349345217956, 0.8653019218632017]
19 Steps Per Batch 64
20 Target KL
21 Track Stats True
22 Update Epochs 8
23 Use Torso True
24 Value Coef 0.609288
Appendix E Additional Results

### v2

@@ 5.3. Theoretical Questions
In two-player zero-sum games, CFR approaches a Nash equilibrium. However, auctions are not zero-sum, and existing guarantees for general-sum games are far weaker. In this case, Morrill et al. (2021) show that CFR is only guaranteed to converge to a counterfactual coarse correlated equilibrum, a much weaker solution concept. Despite this lack of theoretical justification, on our auctions, we found that MCCFR consistently converged to pure Bayes–Nash (near-)equilibrium. Is there a class of games (larger than two-player zero-sum) in which existing MARL algorithms converge to Nash equilibria? Understanding the space of such games would help with selecting an appropriate MARL algorithm.
@@
Acknowledgements.
Thanks to Marc Lanctot and Betty Shea for their technical contributions on early iterations of this work, and to Costa Huang, Paul Milgrom, Julien Perolat, Samuel Sokota, Andrew Vogt, and Albert Zuo for helpful discussions. This work was funded by an NSERC CGS-D scholarship, an NSERC Discovery Grant, a DND/NSERC Discovery Grant Supplement, a CIFAR Canada AI Research Chair (Alberta Machine Intelligence Institute), a Digital Research Alliance of Canada RAC Allocation, awards from Facebook Research and Amazon Research, DARPA award FA8750-19-2-0222, CFDA #12.910 (Air Force Research Laboratory), and through computational resources and services provided by Advanced Research Computing at the University of British Columbia. This work was done in part while several authors were visiting the Simons Institute for the Theory of Computing.
@@ Appendix A Computational Environment
Our experiments were performed on two different compute clusters. Our first cluster consisted of nodes equipped with 32 2.10GHz Intel Xeon E5-2683 v4 CPUs with 40960 KB cache and 96 GB RAM. Our second cluster consisted of 16-core machines with Intel Silver 4216 Cascade Lake 2.1GHz processors and 96 GB RAM. Each run had access to 4 CPUs and 20 GB of RAM.
@@ Appendix C Value Sampling Details
Our game sampler rejects games that we deem strategically uninteresting. We require that every bidder, in all type realizations, is allocated at least one license under the utility-maximizing allocation at opening prices. We solve for the utility-maximizing allocation using a MIP. A bidder that wins nothing is hopelessly weak, slowing down simulations for no added benefit.
@@
Some criteria ensured that we generated games that we were capable of solving. First, as a proxy measure for the length of the auction, we computed the number of rounds it would take for the auction to conclude if each bidder bid for its profit-maximizing bundle at start-of-round prices in every round. We rejected auctions longer than 20 such rounds. Second, we reran the above simulation, only incrementing the price on a single product randomly chosen among those that were overdemanded in each round. This is a crude model of how long the auction could last in the worst case. We rejected simulations that took on average longer than 25 rounds to conclude. Lastly, for 2-player games, we required MCCFR to be able to run at least 25 iterations per second. These checks were all based on the drop-by-player rules.
@@
We made several small changes to the value model for the 3-player games in Section 4.3. For MRVM parameters, we sampled market shares uniformly from 20% to 30% and values per subscriber uniformly from \$35 million to \$45 million. The keypoints on their sigmoidal value function were placed at $\pm$10from their market share. We only required MCCFR run at 1 iteration per second rather than 25.
@@
Appendix D PPO Features
Table 1 lists the features describing an information state that were provided to each PPO agent. All features were normalized such that typical values ranged from 0 to 1.
Feature Description # Variables
Bundle Quantities Number of licenses included in each bundle. $P\times B$
Bundle Activity Total activity points of each bundle. $B$
Values Bidder’s value for each bundle. $B$
Prices History of prices for each bundle. $T\times B$
Submitted Demands This bidder’s bid in each round, one-hot encoded. $T\times B$
Processed Demands This bidder’s bid in each round, after applying bid processing. $T\times B$
Aggregate Demands Sum of all bidders’ processed demands in each round. $T\times P$
Current Activity Total activity points of this bidder’s last processed bundle. 1
Current Exposure Start-of-round cost of this bidder’s last processed bundle. $1$
Table 1. Features provided to each PPO agent. In the # Variables column, $P$ denotes the number of distinct products; $B$ denotes the total number of bundles; and $T$ denotes the maximum length of the auction.
Appendix E PPO Hyperparameters
@@
config[’optimizer_kwargs’] = {’betas’: [beta1, beta2]}
@@
config[’agent_fn_kwargs’] = dict()
@@
config[’agent_fn_kwargs’][’activation’] = str(np.random.choice(
[’relu’, ’tanh’]
config[’agent_fn_kwargs’][’hidden_sizes’] = random.choice(
[[32, 32, 32], [64, 64, 64], [128, 128, 128]
config[’agent_fn_kwargs’][’add_skip_connections’] = bool(np.random.choice(
[True, False]
config[’agent_fn_kwargs’][’use_torso’] = bool(np.random.choice([True, False]))
@@
config[’agent_fn_kwargs’][’actor_hidden_sizes’] = hidden_sizes
config[’agent_fn_kwargs’][’critic_hidden_sizes’] = hidden_sizes
config[’agent_fn_kwargs’][’actor_activation’] = activation
config[’agent_fn_kwargs’][’critic_activation’] = activation
Our final PPO configuration is described in Table 2, and the performance of 60 randomly sampled configurations is plotted in Figure 7.
@@
Activation relu
Add Skip Connections True
Anneal LR True
Architecture AuctionNet
Clip Vloss False
Entropy Coef $3.0\times 10^{-5}$
Gae False
Gamma 1.0
Hidden Sizes [32, 32, 32]
Learning Rate $9\times 10^{-5}$
Max Grad Norm 0.526
Normalize Advantages False
Num Envs 16
Num Minibatches 8
Optimizer adam
Optimizer Betas [0.803, 0.865]
Steps Per Batch 64
Track Stats True
Update Epochs 8
Use Torso True
Value Coef 0.609
Table 2. Our best-performing hyperparameters for PPO.
Appendix F Additional Results
In Figures 8 and 9, we show auction outcomes on the 2-player games, separating equilibria found by each of the two MARL algorithms. Interestingly, we note that there exist some equilibria that are only found by one of the two algorithms. Figure 10 shows auction outcomes on these games assuming bidders follow a straightforward bidding heuristic. Finally, Figure 11 shows the distribution of NashConv for PPO runs on the 2-player games.
@@ Appendix F Additional Results
Figure 11. Distribution of NashConv for PPO on 3-player games, including only runs for which NashConv could be computed in under 1 hour.